40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6087     Entry time: Thu Dec 8 01:26:31 2011
Author: Zach 
Type: Update 
Category: RF System 
Subject: EOM temp sensor modified 

I have modified the EOM temperature sensor circuit for the temperature vs. RAM long-term measurements. The only real change is that the sensor is a 100-kOhm thermistor, instead of a 100-Ohm RTD. These semiconductor thermistors (DigiKey P/N 317-1377-ND) are highly nonlinear and can be much more responsive than RTDs, but this difference is much more noticeable at low temperatures.

Frank had told me that the fractional response of the thermistors was so much higher that I could scale the bridge drive current down by the same factor as the resistance was increasing (i.e., 1 mA -> 1 uA, commensurate with 100 ohms -> 100k) and still see a marked improvement. It turns out that at room temperature for this particular sensor the gain enhancement would only be about ~10x, so I only reduced the drive current to ~10 uA, by INCREASING the drive voltage from ~0.1 V -> 1 V, improving the enhancement to ~100x.

Below is a plot of the real nonlinear response of the thermistor, along with a linear approximation at 298.15 K, which gives a coefficient of ~ -4.67 kOhms/K. The differential bridge output voltage response for the new resistance and current is ~7.5 uV/Ohm 2.5 uV/Ohm, bringing the total temperature response before amplification to ~35 mV/K 11.6 mV/K. Looking at a trend of the FSS_RMTEMP channel over a month, we saw that the maximum PSL table temperature fluctuations were ~2 Kpp, so we aimed the maximize resolution by matching +/- 2 K with +/-10 V at the ADC. This was done by using a gain of ~300 in the AD620 that amplifies the differential bridge output. We found that a gain of ~300 put it pretty close, so the grand total calibration ~ 10.5 V/K 3.5 V/K.

Edit (ZK): I screwed up with calculating the bridge response by a factor of three somehow, so I have stricken and restated the calibrations above

 

 

100k_thermistor_response.png

I took a look at the recently acquired temperature data alongside the RAMmon 11 and 55 signals, and it appears that we're seeing the same sort of fringing effects we usually see, with oscillatory RAM levels for a monotonic change in EOM temperature.The odd bit towards the end is caused by the MC losing lock. 

It is going to be very interesting to find out what causes this fringing.

RAM_with_temp.png

ELOG V3.1.3-